OpenStudy (anonymous):

You would like to have $55,000 in 12 years by making regular, end of year deposits in an annuity that pays 8% compounded annually. How much should you deposit each year? Do not round until the final answer. Then, round to the nearest dollar.

OpenStudy (anonymous):

$2643 per year.. correct?

OpenStudy (anonymous):

A $5,171 B $2,899 C $4,472 D $3,305

OpenStudy (anonymous):

b 55000= x ((1+0.08)^12-1)/0.08

OpenStudy (anonymous):

So the answer would be B ? which formula did you use

OpenStudy (anonymous):

FV=C((1+r)^t −1)/r FV = future value r = rate t = time

OpenStudy (anonymous):

@nphuongsun93 Thank you so much for the help and sorry if i bothered u with all my questions

OpenStudy (anonymous):

lol dw